# Tracing Quenching in Nearby Galaxies Through Inner Surface Mass Density and Cold Gas Content

### v1

@@ Abstract
software: SciPy (Virtanen et al. 2020), astropy (Astropy Collaboration et al. 2013, 2018, 2022), pandas (pandas development team 2020), Matplotlib (Hunter 2007), seaborn (Waskom 2021), numpy (Harris et al. 2020), lifelines (Davidson-Pilon 2024)facilities: Arecibo, GALEX, WISE, IRAM-30m, APEX, SDSS, APO:2.5m
@@ 1 Introduction
Star formation quenching and the processes that regulate it are integral to understanding galaxy evolution. Two important markers of quenching are variations in galaxy structure and a decline in the availability of cold gas. The relationship between cold gas content and galaxy structure has been studied extensively, beginning with early works which established a connection between Hubble morphological type and global cold gas content. The HI gas fraction $(f_{\mathrm{HI}}\equiv M_{\mathrm{HI}}/M_{\ast})$} increases along the Hubble sequence from early- to late-type galaxies (e.g. Roberts 1969; Young & Knezek 1989; Sage 1993; Conselice 2014). There exists less variation in the H2 gas fraction $(f_{\mathrm{H2}}\equiv M_{\mathrm{H2}}/M_{\ast})$ among spiral morphological types, but elliptical types have appreciably lower fractional H2 mass content and often constitute non-detections in targeted surveys (e.g. Young & Scoville 1991; Sage 1993; Roberts & Haynes 1994; Boselli et al. 2014).
@@
Scaling relations between galactic structure and cold gas mass offer a means of quantifying the interplay between gas content and galaxy evolution. There have been many studies of the relationship between HI and H2 content and structural parameters, such as stellar mass surface density (e.g. Saintonge et al. 2011; Catinella et al. 2013; Brown et al. 2015; Saintonge et al. 2017; Catinella et al. 2018; Baker et al. 2021); the half-light or effective radius (e.g. Yesuf & Ho 2019; Lin et al. 2020); Sérsic index (e.g. Fisher et al. 2013; Namiki et al. 2021); and non-parametric measures such as concentration index, asymmetry, and the Gini coefficient (e.g. Catinella et al. 2010; Wang et al. 2011; Kauffmann et al. 2012; Zhou et al. 2018; Namiki et al. 2021; Davis et al. 2022).
@@
The strongest cold gas scaling relations with galactic structure typically involve stellar mass and stellar mass surface density, defined in this work as $\mu_{\ast}\equiv 0.5M_{\ast}$$\pi R_{50,z}^{2}$, where $R_{50,z}$ is the half-light radius computed from the SDSS z band flux. Stellar mass surface density has been used to quantify the prominence of the central bulge component, whereby galaxies become bulge-dominated at $\log\mu_{\ast}>8.7\ M_{\odot}\ \mathrm{kpc}^{-2}$ (Catinella et al. 2010). For nearby galaxies, $\mu_{\ast}$ is anti-correlated with both $f_{\mathrm{HI}}$ and $f_{\mathrm{H2}}$, such that more bulge-dominated galaxies have lower gas fractions (e.g. Catinella et al. 2010; Saintonge et al. 2017). Stellar mass is strongly anti-correlated with atomic gas fraction; $f_{\mathrm{HI}}$ scales with $\log M_{\ast}$ with a slope ranging from 0.3-0.3 to 0.9-0.9 dex, corresponding to a drop of $\sim 70-80\%$ in the gas fraction per dex in $M_{\ast}$ (e.g. Brown et al. 2015; Catinella et al. 2018; Hunt et al. 2020). In contrast, molecular gas fraction shows a weak dependence with $M_{\ast}$, with previous studies finding a near-flat relation, e.g. a median of $\log f_{\mathrm{H_{2}}}\simeq-0.99\pm 0.22$ and a slope consistent with zero over a broad mass range (Saintonge et al. 2011, 2017).
@@
The focus of this paper is to examine the relationship between cold gas content and the stellar mass surface density within 1 kpc of the galactic center ($\Sigma_{1}$) in nearby, massive galaxies. While similar to $\mu_{\ast}$, $\Sigma_{1}$ measures the stellar mass surface density within a smaller, constant radius. $\Sigma_{1}$ and similar parameters (such as the inner surface mass density within 1.5 or 2 kpc) have become popular metrics in recent years to identify and examine galaxies that have been quenched. Early works presented $\Sigma_{1}$ as a proxy for central bulge growth and for central black hole mass ($M_{\mathrm{BH}}$), and have proposed that once galaxies reach a mass-dependent $\Sigma_{1}$ threshold value, their star formation is shut down (Cheung et al. 2012; Fang et al. 2013; Barro et al. 2017). For this reason, $\Sigma_{1}$ is often used as a tracer of galactic evolution from the star-forming to the quiescent phase. As $\Sigma_{1}$ grows at fixed mass, galaxies transition through the green valley from the blue sequence to the red (Fang et al. 2013), and at fixed redshift, galaxies evolve along the star formation main sequence (MS) via inside-out growth, traced by increasing $\Sigma_{1}$ (Barro et al. 2017).
@@
$\Sigma_{1}$ is tightly correlated with $M_{\ast}$ (e.g. Fang et al. 2013). The mass-trend-removed parameter $\Delta\Sigma_{1}$ is defined relative to the structural valley within the $\Sigma_{1}-M_{\ast}$ plane. This is a visual feature in the plot which, at fixed mass, appears as a trough between two higher density populations of low and high $\Sigma_{1}$ (Luo et al. 2020). While not exactly the same as the star-formation green valley, Luo et al. (2020) showed that $\Delta\Sigma_{1}$ as defined in this manner is related to the star-forming/quiescent division in massive galaxies. Luo et al. (2020) calculated a parabolic fit to mark the structural valley in the $\Sigma_{1}-M_{\ast}$ plane, and $\Delta\Sigma_{1}$ is defined as the offset in $\log\Sigma_{1}$ from this line at fixed mass. With the mass trend removed, $\Delta\Sigma_{1}$ provides a more robust probe of the relationship between inner stellar mass surface density growth and global cold gas fraction.
@@
When $\Sigma_{1}$ or $\Delta\Sigma_{1}$ is plotted against SFR, or alternatively against other parameters that trace SFR–such as color, specific SFR (sSFR), or distance from the MS–a distinctive ‘elbow’ or ‘L-shaped’ pattern emerges in the distribution of galaxies in bins of $M_{\ast}$ (Fang et al. 2013; Barro et al. 2017; Lee et al. 2018; Luo et al. 2020; Guo et al. 2021). It has been proposed that there is more than one possible track along which galaxies evolve as they transition from star-forming to quiescent. For example, in the plane of $\Sigma_{1}$ vs. sSFR, evolution may be characterized by either gas-rich or gas-poor quenching processes, which correspond to quenching at earlier and later times, respectively (Barro et al. 2013, 2017). For satellites, differences in evolution within this plane may involve quenching via bulge growth driven by central star formation, or through global mass-loss processes like tidal stripping (Woo et al. 2016). In any case, the consensus of the $\Sigma_{1}$ literature is that galaxies must build up their central bulge as a prerequisite for quenching.
@@
The proposed mechanisms of massive galaxy quenching are too complex to discuss in depth here, and vary according to a variety of galaxy properties. Theoretical and simulation analyses contend that AGN feedback is a prominent – even necessary – mechanism for massive galaxy quenching (e.g. Croton et al. 2006; Dubois et al. 2013; Su et al. 2019; Piotrowska et al. 2021). Because of the close relationship between $\Sigma_{1}$ and the growth of the central black hole, we focus on the role of AGN feedback in gas removal and quenching and its connection to $\Delta\Sigma_{1}$.
@@
It has been established that $f_{\mathrm{H2}}$ is tightly correlated with sSFR (Saintonge et al. 2017). In the plane of sSFR and $\Delta\Sigma_{1}$, galaxies follow a bimodal distribution which approximately separates star-forming from quiescent galaxies (Luo et al. 2020). These two relationships motivate us to investigate whether central bulge growth is also related to global cold gas content, particularly H2. Moreover, the scaling of cold gas content with $\Sigma_{1}$ has not been thoroughly explored in previous research. A few studies have analyzed the relationship between $\Sigma_{1}$ and cold gas content within the framework of cosmological simulations. Ma et al. (2022) and Khan et al. (2025) each concluded that $\Sigma_{1}$ is connected to a decline in cold gas in galaxy centers within the IllustrisTNG and NIHAO simulations, respectively, and offered AGN feedback as a plausible explanation. In this work, we will use observational data to compare the relationship between $\Sigma_{1}$ and global cold gas content to established gas scaling relations with $M_{\ast}$ and $\mu_{\ast}$. With the availability of molecular and atomic cold gas catalogs for representative samples, we are also positioned to explore the usefulness of $\Sigma_{1}$ and $\Delta\Sigma_{1}$ as indicators of galaxy quenching through an analysis of the cold gas reservoirs of galaxies with a range of $\Delta\Sigma_{1}$ values.
@@
2 Sample Selection
The Extended GALEX Arecibo SDSS Survey (xGASS; Catinella et al. 2018) and Extended CO Legacy Database for GASS (xCOLDGASS; Saintonge et al. 2017) provide publicly available HI and H2 catalogs for mass-selected samples of nearby galaxies. Spectral measurements were obtained from the MPA/JHU DR7 catalog [^1] https://wwwmpa.mpa-garching.mpg.de/SDSS/DR7/ for the purposes of BPT classification (Brinchmann et al. 2004). The samples were then combined with $\Sigma_{1}$ data (Luo et al. 2020).
[…]
@@ 2 Sample Selection
Cross-Matched with $\Sigma_{1}$ Catalog (Luo et al. 2020) $448$
@@
Cross-Matched with $\Sigma_{1}$ Catalog (Luo et al. 2020) $226$
@@
Figure 1: (Left) The contours represent the SDSS main sample on the SFR-$M_{\ast}$ plane, and we include the line from Woo et al. 2015 to indicate objects that are either star-forming (above the line) or non-star-forming (below the line). The points comprise our final HI sample from the original xGASS sample. Code 2 xGASS detections are square markers to distinguish them as “less reliable” or “marginal” detections. (Right) The contours of the full SDSS sample are now overlaid with the points comprising our H2 sample, which was derived from the full xCOLDGASS sample. In both plots, non-detections are indicated by triangle markers. For further explanation, see Sections 2.2 and 2.3.
2.1 $\Sigma_{1}$ Data
Luo et al. (2020) measured $\Sigma_{1}$ values for a volume- and mass-limited sample of >40,000>40,000 SDSS galaxies. In particular, they focused on central galaxies with good viewing geometries, imposing the following restrictions: stellar mass $M_{\ast}>10^{9.5}M_{\odot}$; axis ratio $b/a>0.5$; Sérsic index $0.5<n<6$; and redshift $0.02<z<0.07$. See Luo et al. (2020) for details of their survey methods.
@@ 2.1 $\Sigma_{1}$ Data
Since we are using the $\Sigma_{1}$ values calculated by Luo et al. (2020), we use the same stellar mass data for consistency, which are obtained from the MPA/JHU DR7 value-added catalog. We also adopt their computed $\Delta\Sigma_{1}$ values in our analysis.
2.2 HI sample
xGASS is a targeted survey of galaxies with $10^{9}M_{\odot}<M_{\ast}<10^{11.5}M_{\odot}$ and redshifts $0.01<z<0.05$. Integration times were set by the requirement that certain upper limits be reached unless a measurement could be obtained with a sufficient S/N ratio, allowing for greater sensitivity in the low $f_{\mathrm{HI}}$ regime. The distribution of galaxies is approximately uniform in stellar mass bins between $\log M_{\ast}/M_{\odot}=10-11$, and decreases steadily for $\log M_{\ast}/M_{\odot}>11$. The representative sample contains 1179 galaxies, a fraction of which are galaxies with archival HI data obtained from ALFALFA or the Cornell HI Archive, which were not remeasured for survey efficiency. For further information on how HI data were collected and chosen for the xGASS sample, see Catinella et al. (2018), and for general GASS survey methods, see Catinella et al. (2010) and Catinella et al. (2013).
@@ 2.2 HI sample
2.3 H2 sample
We extract our H2 data from the xCOLDGASS sample, which spans a mass of $10^{9}<M_{\ast}/M_{\odot}<10^{11.5}$ and redshift range of $0.01<z<0.05$ (Saintonge et al. 2017). The methods of xCOLDGASS are complementary to those of xGASS, and there are 477 galaxies common to both samples.
In xCOLDGASS, H2 masses are extrapolated from CO(1-0) emission line flux measurements, which serve as a tracer for molecular gas. For details about the sampling and detection methods of xCOLDGASS, see Saintonge et al. (2017). Non-detections are reported with $3\sigma$ upper limits in the xCOLDGASS tables, and all detections have S/N >5>5 by survey design. Within our sample, there are 118 detections and 83 non-detections.
@@ 2.4 Our Samples in the BPT Diagram
[…]
Galaxies located above the Kewley et al. (2001) line are classified as AGN candidates. Our HI sample contains 163 such candidates ($48\%$ of the sample) and our H2 sample contains 85 ($42\%$). We note that AGN are somewhat overrepresented in our sample compared to the Luo et al. (2020) sample, in which $\sim 35\%$ of galaxies meeting their selection criteria are AGN for $M_{\ast}>10^{10}M_{\odot}$ and $0.02\leq z\leq 0.05$. This is discussed further in Section 3.4.
Figure 3: (Left) The BPT diagram for our HI sample, and (right) for our H2 sample. We show the theoretical relation of Kewley et al. (2001) and empirical relation of Kauffmann et al. (2003), which separate galaxies based on their dominant source of gas ionization.

### v2

@@ Abstract
software: SciPy (SciPy20), astropy (astropy13; astropy18; astropy22), pandas (pandas20), Matplotlib (Hunter07), seaborn (Waskom21), numpy (Harris20), lifelines (davidson_pilon_2024_14007206)facilities: Arecibo, GALEX, WISE, IRAM-30m, APEX, SDSS, APO:2.5m
@@ I Introduction
Star formation quenching and the processes that regulate it are integral to understanding galaxy evolution. Two important markers of quenching are variations in galaxy structure and a decline in the availability of cold gas. The relationship between cold gas content and galaxy structure has been studied extensively, beginning with early works which established a connection between Hubble morphological type and global cold gas content. The HI gas fraction $(f_{\mathrm{HI}}\equiv M_{\mathrm{HI}}/M_{\ast})$} increases along the Hubble sequence from early- to late-type galaxies (e.g. Roberts69; Young89; Sage93; Conselice14). There exists less variation in the H2 gas fraction $(f_{\mathrm{H2}}\equiv M_{\mathrm{H2}}/M_{\ast})$ among spiral morphological types, but elliptical types have appreciably lower fractional H2 mass content and often constitute non-detections in targeted surveys (e.g. Young91; Sage93; Roberts94; Boselli14).
@@
Scaling relations between galactic structure and cold gas mass offer a means of quantifying the interplay between gas content and galaxy evolution. There have been many studies of the relationship between HI and H2 content and structural parameters, such as stellar mass surface density (e.g. Saintonge11; Catinella13; Brown15; Saintonge17; Catinella18; Baker21); the half-light or effective radius (e.g. Yesuf19; Lin20); Sérsic index (e.g. Fisher13; Namiki21); and non-parametric measures such as concentration index, asymmetry, and the Gini coefficient (e.g. Catinella10; Wang11; Kauffmann12; Zhou18; Namiki21; Davis22).
@@
The strongest cold gas scaling relations with galactic structure typically involve stellar mass and stellar mass surface density, defined in this work as $\mu_{\ast}\equiv 0.5M_{\ast}$$\pi R_{50,z}^{2}$, where $R_{50,z}$ is the half-light radius computed from the SDSS z band flux. Stellar mass surface density has been used to quantify the prominence of the central bulge component, whereby galaxies become bulge-dominated at $\log\mu_{\ast}>8.7\ M_{\odot}\ \mathrm{kpc}^{-2}$ (Catinella10). For nearby galaxies, $\mu_{\ast}$ is anti-correlated with both $f_{\mathrm{HI}}$ and $f_{\mathrm{H2}}$, such that more bulge-dominated galaxies have lower gas fractions (e.g. Catinella10; Saintonge17). Stellar mass is strongly anti-correlated with atomic gas fraction; $f_{\mathrm{HI}}$ scales with $\log M_{\ast}$ with a slope ranging from 0.3-0.3 to 0.9-0.9 dex, corresponding to a drop of $\sim 70-80\%$ in the gas fraction per dex in $M_{\ast}$ (e.g. Brown15; Catinella18; Hunt20). In contrast, molecular gas fraction shows a weak dependence with $M_{\ast}$, with previous studies finding a near-flat relation, e.g. a median of $\log f_{\mathrm{H_{2}}}\simeq-0.99\pm 0.22$ and a slope consistent with zero over a broad mass range (Saintonge11; Saintonge17).
@@
The focus of this paper is to examine the relationship between cold gas content and the stellar mass surface density within 1 kpc of the galactic center ($\Sigma_{1}$) in nearby, massive galaxies. While similar to $\mu_{\ast}$, $\Sigma_{1}$ measures the stellar mass surface density within a smaller, constant radius. $\Sigma_{1}$ and similar parameters (such as the inner surface mass density within 1.5 or 2 kpc) have become popular metrics in recent years to identify and examine galaxies that have been quenched. Early works presented $\Sigma_{1}$ as a proxy for central bulge growth and for central black hole mass ($M_{\mathrm{BH}}$), and have proposed that once galaxies reach a mass-dependent $\Sigma_{1}$ threshold value, their star formation is shut down (Cheung12; Fang13; Barro17). For this reason, $\Sigma_{1}$ is often used as a tracer of galactic evolution from the star-forming to the quiescent phase. As $\Sigma_{1}$ grows at fixed mass, galaxies transition through the green valley from the blue sequence to the red (Fang13), and at fixed redshift, galaxies evolve along the star formation main sequence (MS) via inside-out growth, traced by increasing $\Sigma_{1}$ (Barro17).
@@
$\Sigma_{1}$ is tightly correlated with $M_{\ast}$ (e.g. Fang13). The mass-trend-removed parameter $\Delta\Sigma_{1}$ is defined relative to the structural valley within the $\Sigma_{1}-M_{\ast}$ plane. This is a visual feature in the plot which, at fixed mass, appears as a trough between two higher density populations of low and high $\Sigma_{1}$ (Luo20). While not exactly the same as the star-formation green valley, Luo20 showed that $\Delta\Sigma_{1}$ as defined in this manner is related to the star-forming/quiescent division in massive galaxies. Luo20 calculated a parabolic fit to mark the structural valley in the $\Sigma_{1}-M_{\ast}$ plane, and $\Delta\Sigma_{1}$ is defined as the offset in $\log\Sigma_{1}$ from this line at fixed mass. With the mass trend removed, $\Delta\Sigma_{1}$ provides a more robust probe of the relationship between inner stellar mass surface density growth and global cold gas fraction.
@@
When $\Sigma_{1}$ or $\Delta\Sigma_{1}$ is plotted against SFR, or alternatively against other parameters that trace SFR–such as color, specific SFR (sSFR), or distance from the MS–a distinctive ‘elbow’ or ‘L-shaped’ pattern emerges in the distribution of galaxies in bins of $M_{\ast}$ (Fang13; Barro17; Lee2018; Luo20; Guo2021). It has been proposed that there is more than one possible track along which galaxies evolve as they transition from star-forming to quiescent. For example, in the plane of $\Sigma_{1}$ vs. sSFR, evolution may be characterized by either gas-rich or gas-poor quenching processes, which correspond to quenching at earlier and later times, respectively (Barro13; Barro17). For satellites, differences in evolution within this plane may involve quenching via bulge growth driven by central star formation, or through global mass-loss processes like tidal stripping (Woo16). In any case, the consensus of the $\Sigma_{1}$ literature is that galaxies must build up their central bulge as a prerequisite for quenching.
@@
The proposed mechanisms of massive galaxy quenching are too complex to discuss in depth here, and vary according to a variety of galaxy properties. Theoretical and simulation analyses contend that AGN feedback is a prominent – even necessary – mechanism for massive galaxy quenching (e.g. Croton_2006; Dubois2013; Su2019; Piotrowska2021). Because of the close relationship between $\Sigma_{1}$ and the growth of the central black hole, we focus on the role of AGN feedback in gas removal and quenching and its connection to $\Delta\Sigma_{1}$.
@@
It has been established that $f_{\mathrm{H2}}$ is tightly correlated with sSFR (Saintonge17). In the plane of sSFR and $\Delta\Sigma_{1}$, galaxies follow a bimodal distribution which approximately separates star-forming from quiescent galaxies (Luo20). These two relationships motivate us to investigate whether central bulge growth is also related to global cold gas content, particularly H2. Moreover, the scaling of cold gas content with $\Sigma_{1}$ has not been thoroughly explored in previous research. A few studies have analyzed the relationship between $\Sigma_{1}$ and cold gas content within the framework of cosmological simulations. Ma2022 and Khan_2025 each concluded that $\Sigma_{1}$ is connected to a decline in cold gas in galaxy centers within the IllustrisTNG and NIHAO simulations, respectively, and offered AGN feedback as a plausible explanation. In this work, we will use observational data to compare the relationship between $\Sigma_{1}$ and global cold gas content to established gas scaling relations with $M_{\ast}$ and $\mu_{\ast}$. With the availability of molecular and atomic cold gas catalogs for representative samples, we are also positioned to explore the usefulness of $\Sigma_{1}$ and $\Delta\Sigma_{1}$ as indicators of galaxy quenching through an analysis of the cold gas reservoirs of galaxies with a range of $\Delta\Sigma_{1}$ values.
@@
II Sample Selection
The Extended GALEX Arecibo SDSS Survey (xGASS; Catinella18) and Extended CO Legacy Database for GASS (xCOLDGASS; Saintonge17) provide publicly available HI and H2 catalogs for mass-selected samples of nearby galaxies. Spectral measurements were obtained from the MPA/JHU DR7 catalog [^1] https://wwwmpa.mpa-garching.mpg.de/SDSS/DR7/ for the purposes of BPT classification (Brinchmann04). The samples were then combined with $\Sigma_{1}$ data (Luo20).
In Figure 1, we show the distribution of our HI and H2 samples in the SFR-$M_{\ast}$ plane, compared to the SDSS DR7 Catalog (Abazajian09). To identify galaxies as either star-forming or quiescent, we adopted the criterion set by Woo15, who divided SDSS galaxies in the SFR-$M_{\ast}$ plane according to the line $\log\text{SFR}=0.74\log M_{\ast}-8.22$. Galaxies above this line are classified as star-forming, and those below it are considered quiescent. This line also fits our data well, even though the SFR values we use originate from a different source (the xGASS and xCOLDGASS catalogs).
@@ II Sample Selection
Cross-Matched with $\Sigma_{1}$ Catalog (Luo20) $448$
@@
Cross-Matched with $\Sigma_{1}$ Catalog (Luo20) $226$
@@
Figure 1: (Left) The contours represent the SDSS main sample on the SFR-$M_{\ast}$ plane, and we include the line from Woo15 to indicate objects that are either star-forming (above the line) or non-star-forming (below the line). The points comprise our final HI sample from the original xGASS sample. Code 2 xGASS detections are square markers to distinguish them as “less reliable” or “marginal” detections. (Right) The contours of the full SDSS sample are now overlaid with the points comprising our H2 sample, which was derived from the full xCOLDGASS sample. In both plots, non-detections are indicated by triangle markers. For further explanation, see Sections II.2 and II.3.
II.1 $\Sigma_{1}$ Data
Luo20 measured $\Sigma_{1}$ values for a volume- and mass-limited sample of >40,000>40,000 SDSS galaxies. In particular, they focused on central galaxies with good viewing geometries, imposing the following restrictions: stellar mass $M_{\ast}>10^{9.5}M_{\odot}$; axis ratio $b/a>0.5$; Sérsic index $0.5<n<6$; and redshift $0.02<z<0.07$. See Luo20 for details of their survey methods.
@@ II.1 $\Sigma_{1}$ Data
Since we are using the $\Sigma_{1}$ values calculated by Luo20, we use the same stellar mass data for consistency, which are obtained from the MPA/JHU DR7 value-added catalog. We also adopt their computed $\Delta\Sigma_{1}$ values in our analysis.
II.2 HI sample
xGASS is a targeted survey of galaxies with $10^{9}M_{\odot}<M_{\ast}<10^{11.5}M_{\odot}$ and redshifts $0.01<z<0.05$. Integration times were set by the requirement that certain upper limits be reached unless a measurement could be obtained with a sufficient S/N ratio, allowing for greater sensitivity in the low $f_{\mathrm{HI}}$ regime. The distribution of galaxies is approximately uniform in stellar mass bins between $\log M_{\ast}/M_{\odot}=10-11$, and decreases steadily for $\log M_{\ast}/M_{\odot}>11$. The representative sample contains 1179 galaxies, a fraction of which are galaxies with archival HI data obtained from ALFALFA or the Cornell HI Archive, which were not remeasured for survey efficiency. For further information on how HI data were collected and chosen for the xGASS sample, see Catinella18, and for general GASS survey methods, see Catinella10 and Catinella13.
@@ II.2 HI sample
II.3 H2 sample
We extract our H2 data from the xCOLDGASS sample, which spans a mass of $10^{9}<M_{\ast}/M_{\odot}<10^{11.5}$ and redshift range of $0.01<z<0.05$ (Saintonge17). The methods of xCOLDGASS are complementary to those of xGASS, and there are 477 galaxies common to both samples.
In xCOLDGASS, H2 masses are extrapolated from CO(1-0) emission line flux measurements, which serve as a tracer for molecular gas. For details about the sampling and detection methods of xCOLDGASS, see Saintonge17. Non-detections are reported with $3\sigma$ upper limits in the xCOLDGASS tables, and all detections have S/N >5>5 by survey design. Within our sample, there are 118 detections and 83 non-detections.
@@ II.4 Our Samples in the BPT Diagram
Figure 3 shows the distribution of our sample in the BPT diagram. There are fewer star-forming galaxies, demarcated by their location below the Kauffmann03 line, than typically observed for larger, more representative samples. There is a documented mass trend within the BPT diagram whereby the star-forming branch is dominated by lower-mass galaxies (e.g. Kumari21); since our sample is limited to higher-mass objects ($\log M_{\ast}/M_{\odot}\gtrsim 10$), the distribution of star-forming galaxies is lacking, particularly for $\log(\text{[NII]/H}\alpha)<-0.6$.
Galaxies located above the Kewley01 line are classified as AGN candidates. Our HI sample contains 163 such candidates ($48\%$ of the sample) and our H2 sample contains 85 ($42\%$). We note that AGN are somewhat overrepresented in our sample compared to the Luo20 sample, in which $\sim 35\%$ of galaxies meeting their selection criteria are AGN for $M_{\ast}>10^{10}M_{\odot}$ and $0.02\leq z\leq 0.05$. This is discussed further in Section III.4.
Figure 3: (Left) The BPT diagram for our HI sample, and (right) for our H2 sample. We show the theoretical relation of Kewley01 and empirical relation of Kauffmann03, which separate galaxies based on their dominant source of gas ionization.